Mary walked 4 laps around the park track. It took Mary 4 minutes to walk the first
lap. The second lap took 25% longer than the first lap. The third lap took 20%
longer than the second lap. The final lap took 50% longer than the third. What was
Mary�s total walking time for the 4 laps?

Found 3 solutions by mananth, ikleyn, josgarithmetic:
Answer by mananth(16949)

Mary walked 4 laps around the park track. It took Mary 4 minutes to walk the first
lap. The second lap took 25% longer than the first lap. The third lap took 20%
longer than the second lap. The final lap took 50% longer than the third. What was
Mary’s total walking time for the 4 laps?
First lap is 4 mins.
Second lap = 25/100*4 +4 =5
Third lap= 20/100*5 + 5 =6
Fourth lap = 50/100*6 +6= 9
The total time taken for all the laps = 5+6+9= 20 mins
M ananth ---------- mananth@hotmail.com

Answer by ikleyn(53875)

.
Mary walked 4 laps around the park track. It took Mary 4 minutes to walk the first
lap. The second lap took 25% longer than the first lap. The third lap took 20%
longer than the second lap. The final lap took 50% longer than the third. What was
Mary�s total walking time for the 4 laps?
~~~~~~~~~~~~~~~~~~~~~


The correct answer to this problem is not 5+6+9 = 20 minutes, as it is stated in the @mananth solution.

The correct answer is 4 + 5 + 6 + 9 = 24 minutes.



Answer by josgarithmetic(39823)

LAP        TIME
 1          4
 2          4+1=5
 3          5+1=6
 4           6+3=9

Total time, 4+5+6+9